OpenStudy (anonymous):

3x+19/4 - -x-37/5=15

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

Are you still typing??

OpenStudy (radar):

3x + 19/4 - -x -37/5 -------------- 4x + 243/20 =300/20 4x=57/20 x=(57/20)(1/4)=57/80

OpenStudy (radar):

Should the answer be in decimal or a fraction?

OpenStudy (anonymous):

Fraction

OpenStudy (anonymous):

How did you get 243?

OpenStudy (anonymous):

Help mE?

OpenStudy (anonymous):

3x+19/4 - -x-37/5=15 collect like terms 3x+x+19/4-37/5=15 4x=15-19/4+37/5 You have a common denominator of 20, so multply everything by 20 20(4x)=20*15-20*(19/4)+20(37/5) 80x=300-95+148 = 353 divide by 80 x=353/80

OpenStudy (radar):

Now we have two answers provided. x=57/80 and x=353/80 Sooooo which one is the correct solution, since I got the x=57/80, I feel bound to verify its accuracy. How do I do this? I substitute it into the original problem which I repeat here: 3x + 19/4 --x - 37/5 = 15 Now for the first ambiguity. Is this a subtraction problem where the fraction (-x-37/5) is subtracted from (3x + 19/4) I am assuming, so the problem is: 3x + 19/4 - (-x-37/5) = 15 Substituting for the solution of x=57/80. 3(57/80) + 19/4 + 57/80 + 37/5 =15 171/80 + 19/4 +57/80 + 37/5 = 15 228/80 + 19/4 +37/5 = 15 228/80 + 380/80 + 592/80 = 15 1200/80 = 15 15=15 Checks. The problem could be also looked at like this only the -x is to be subtracted: 3x + 19/4 - -x -37/5 = 15 which can be rewritten as: 3x + 19/4 +x - 37/5 = 15 see the difference. now collect similar terms; 4x +19/4 - 37/5 = 15 4x=15 - 19/4 +37/5 common denominator is 20 4x= 300/20 - 95/20 + 148/20 4x = 353/20 x = 353/80 which is the second solution Only luvstuding can determine what was meant by the problem.